An investor buys 100 shares of ENCANA at $94. OO/share and writes a call with a strike price of $80.00 for $8.00. The positions are closed with the stock is at $90.00. The maximum gross gain on the transactions is:

Select one:

a. $600.00

b. $400.00

c. -$400.00

d. -$600.00
Answer #1

Profit from the shares = (Selling price - Purchasing price) * Number of shares

Profit from the shares = (90 - 94) * 100 = -$400

Profit from selling call option = (-max(St - X, 0) + premium received) * 100

Profit from selling call option = (-max(90 - 80, 0) + 8 ) * 100

Profit from selling call option = (-max(10, 0) + 8) * 100

Profit from selling call option = (-10 + 8) * 100

Profit from selling call option = -2 * 100

Profit from selling call option = -$200

The maximum gross gain = -400 - 200 = -$600

Option d is correct

Can you please upvote? Thank You :-)